OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

(a)+(a+1)+(a+2)=267. Can you take it from here?

OpenStudy (anonymous):

Noooooo. Could you help me a little more?

OpenStudy (anonymous):

well if you add them you will get 3a+3 = 267 or 3a = 264. so a = 264/3. and the largest integer would be a + 2.

OpenStudy (anonymous):

So the final answer is a+2?

OpenStudy (anonymous):

do you know algebra?

OpenStudy (anonymous):

The largest integer would be 90

OpenStudy (anonymous):

Not at all. Im horrible at it.

OpenStudy (anonymous):

How did you get that? @Jesusc

OpenStudy (anonymous):

well you need to plug in the value for a. as a just represents a number. So the final answer would be 264/3 + 2

OpenStudy (anonymous):

and that is 90 like jesusc said.

OpenStudy (anonymous):

So the answer is 90?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you guys.

OpenStudy (anonymous):

JBrandonS explained it very nicely - it's an algebraic question